OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0

OpenStudy (anonymous):

Step 1: Rearrange the two equations in a way where the x&y coefficients are one side while the other numbers (without variables of course) are on the other side. Step 2: It should be something like this: 6x+y = 4 5x+y = 7 Let's say that [ 6 1 ] [ 5 1 ] is equal to |A| Step 3: Set up the determinant equal to the x or y value. [ 6 1 ] _ [ 4 ] [ 5 1 ] - [ 7 ] Step 4: Solve for |A| det(|A|) =det([ 6 1 ] [ 5 1 ]) = 1 Step 5 Replace the first column of the 2 x 2 matrix with [4] [7] In this case, the determinant would be [ 4 1 ] [ 7 1 ] This determinant would be |Ax| = -3. Step 6: Replace the second column of the 2x2 matrix with [ 6 4 ] [ 5 7 ] In this case, the determinant would be (6*7) - (4*5) = 35 - 20 = 15 This determinant would be |Ay| = +15 Step 7: Solving for x and y. x and y could be solved through the following formulas: x = |Ax|/|A| = -3 y = |Ay|/|A| = +15 I hope this helps =D

OpenStudy (anonymous):

You can always plug back the equation to see if the answer is correct or not. I actually did it myself and it ended up working for me.